# The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

### 11 Apr 2026 v2

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) is a practical, scalable way to improve large language models on math, code, and other structured tasks. However, we argue that many headline RLVR gains are not yet well validated because reports often conflate policy improvement with three confounds: (i) budget mismatch between RLVR and baseline evaluation, (ii) attempt inflation and calibration drift that convert abstentions into confident answers, and (iii) data contamination in benchmarks. Using budget-matched reproductions and partial-prompt contamination probes, we find that several widely cited gaps shrink substantially or disappear once budgets, prompts, and dataset versions are matched, and contaminated sets are treated as memorization probes rather than evidence of reasoning. This does not mean that RLVR is ineffective, but it implies that current measurements often overstate capability gains and obscure reliability costs. We therefore propose a compact, tax-aware minimum standard for RLVR training and evaluation: budget-matched saturation curves with variance, calibration, and abstention tracking, one judge robustness stress test when LLM judges are used, and an explicit contamination screen. With these controls, RLVR remains effective and deployable in verifiable domains, but reasoning gains should be treated as provisional without them.
@@
Fang Wu1,*†   Aaron Tu2,*   Weihao Xuan3,*   Heli Qi4,*   Xu Huang5   Qingcheng Zeng6 Shayan Talaei1   Yijia Xiao7   Peng Xia8   Xiangru Tang9   Yuchen Zhuang5   Bing Hu10 Hanqun Cao11   Wenqi Shi12   Rui Yang14   Nan Liu14   Huaxiu Yao8   Ge Liu13 Li Erran Li15   Amin Saberi1   Naoto Yokoya3   Jure Leskovec1   Yejin Choi1,† 1Stanford University  2UC Berkeley  3The University of Tokyo  4Waseda University 5Georgia Tech  6Northwestern University  7UCLA  8UNC Chapel Hill  9Yale University 10Independent Researcher  11CUHK  12UT Southwestern Medical Center  13UIUC 14National University of Singapore  15Amazon AWS AI *Equal Contribution   Corresponding Authors
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a leading post-training route for improving large language models on math and code tasks (Luong et al., 2024; Wen et al., 2025a). By optimizing against automatically computable signals such as unit tests for programs, exact numeric or string matches for math, or retrieval-grounded checks for citations, RLVR promises a path to better performance. Recent work reports large benchmark gains across domains. Appendix Figure 2 shows a trend plot of RLVR-tagged paper activity and AIME-24/25 performance with striking monthly gains. Our position is that RLVR is effective and deployable for verifiable domains, but “reasoning” gains should be treated as provisional without a small core of tax-aware controls.
@@
A central question that frames the current debate is: Does RLVR genuinely impart new reasoning capability, or does it mainly sharpen selection among behaviors the base model already knows how to produce? Budget-parity controlled studies show that base models can narrow or erase RLVR gaps when given matched sampling budgets, consistent with smarter search rather than capability expansion (Yue et al., 2025; Wu et al., 2025c). At the same time, several regimes report gains that are hard to recover by sampling alone: explicitly optimizing the multi-sample objective (pass@k training) (Wang et al., 2025), curriculum via self-play with variational problem synthesis (Zhang et al., 2025a), and distribution-aware reward shaping that counters rank bias and diversity collapse (Liu et al., 2025b), alongside longer-horizon RL schedules and unlikeliness rewards (Liu et al., 2025a; He et al., 2025a). Metric choice can flip conclusions: answer-only scores may diverge from process-aware metrics that require both a correct answer and a valid chain of thought.
@@
A second complication is what we call the RLVR tax: unintended empirical side effects that ride along with apparent gains under current reasoning-style post-training (reduced abstention, miscalibration, instruction-fidelity drift, and a larger safety/privacy surface due to longer traces). This tax is not inevitable and not unique to RLVR; similar patterns arise under reasoning-heavy SFT and RLHF (Wu et al., 2025b). We focus on RLVR because verifiable objectives and open weights make these trade-offs measurable at relatively low cost. In practice, RLVR tends to reduce abstention and increase stated confidence, sometimes even when answers are wrong, thereby shifting risk from “I don’t know” to assertive errors (Song et al., 2025; Mei et al., 2025; Yao et al., 2025). It can also chip away at instruction fidelity on longer generations, where adhering to formats or constraints becomes harder as chains grow (Fu et al., 2025; Li et al., 2025). Finally, longer, more explicit reasoning traces expand the attack and leakage surface, raising jailbreak success and privacy exposure if left unchecked (Zhou et al., 2025a; Jiang et al., 2025; Ackerman and Panickssery, 2025; Green et al., 2025; Huang et al., 2025; Zhang et al., 2025b).
@@
A third complication is measurement. Reported advances are sensitive to sampling budgets and decoding settings, to the stability of LLM-as-a-judge evaluators, to calibration drift, and to data provenance (Hochlehnert et al., 2025; Chandak et al., 2025; Brown and others, 2024; Muennighoff et al., 2025; Zhao et al., 2025; Wang et al., 2024a, b; Leng et al., 2024; Shen et al., 2025; Ye et al., 2026; He et al., 2025b; Mirzadeh et al., 2024; Liu et al., 2024b; Xu et al., 2025; Wu and others, 2024; Wu et al., 2025a, ). When budgets are matched, judges are stress-tested, and datasets are versioned and decontaminated, reported improvement gaps can diminish, suggesting that part of the apparent progress reflects evaluation design rather than durable capability (Hochlehnert et al., 2025; Chandak et al., 2025). We organize the paper around three threads: the RLVR tax, evaluation pitfalls, and data contamination, culminating in a compact tax-aware minimum standard. A visual roadmap is provided in Figure 1.
@@
We use RLVR to mean post-training that optimizes LLMs against automatically checkable signals rather than human preference models (Ouyang and others, 2022). Typical implementations adopt PPO/GRPO/DAPO-style updates with a KL penalty to the base policy, often mixing offline rollouts with online sampling, adding entropy regularization, and using selective filtering or gating. Rewards are frequently componentized (for correctness, grounding/citation sufficiency, and calibrated refusal) and introduced in stages. Our analysis centers on open-weight models fine-tuned for math/code/question answering (QA) with verifiable objectives, in both single- and multi-domain regimes. We deliberately avoid a catalogue of low-level algorithmic variants except where they bear on reliability (e.g., unlikeliness rewards that counter rank bias, or curricula that change variance). The primary threats considered are privacy leakage from long chains of thought (CoT) and increased jailbreak susceptibility during or after RLVR.
@@
Our contributions are threefold: (1) Measurement: we provide budget-matched reproductions and a unified gap table that isolates from budget mismatch, judge and template drift, and dataset versioning (Table 2). (2) Tax: we quantify attempt inflation and the resulting calibration and refusal costs under RLVR and matched SFT controls (Table 1). (3) Protocol: based on these findings, we distill a compact tax-aware minimum standard for RLVR training and evaluation, intended as a minimum credible reporting bar, instantiated in section 6. The paper is organized accordingly: we (i) synthesize evidence on sharpening vs. expansion (section 2); (ii) analyze the RLVR tax and how it distorts reported improvements (section 3); (iii) reproduce representative gaps under parity controls (section 4); (iv) audit contamination with partial-prompt probes (section 5); and (v) present the tax-aware minimum standard (section 6).
@@ 2.1 Skeptical Lens: Sharpening and Sampling Effects
RLVR often appears to improve sample efficiency and steer the model toward high-reward regions already present in the base distribution, rather than broaden fundamental capability (Gandhi et al., 2025; Shah and others, 2025). When base models are evaluated under matched sampling budgets (i.e., pass@k), gaps between base and RLVR-trained models often shrink (Yue et al., 2025), consistent with sharpening rather than expansion. Empirically, outputs can collapse toward a dominant pretraining mode with reduced diversity (Zhao and others, 2025), and theory suggests that after RLVR, support shrinkage tends to outweigh support expansion (Wu et al., 2025c). Stress tests add caution: performance collapses at higher complexities have been reported even with ample compute (Shojaee et al., 2025), although some failures trace to evaluation artifacts such as token limits and unsolvable items, underscoring the need for budget parity, seed control, and dataset hygiene (see section 4(Opus and Lawsen, 2025). Finally, several works separate knowledge injection from policy optimization: pipelines that use SFT or distillation to inject knowledge and RLVR to select among candidates typically beat pure GRPO, suggesting RLVR optimizes within learned support (Ma et al., 2025; Wan et al., 2025; Chen et al., 2025a); guidance and self-distillation further help over vanilla GRPO (Nath et al., 2025; Liu et al., 2025c).
@@ 2.2 Optimistic Lens: Signals of Capability Broadening
Answer-only metrics (e.g., pass@k) can show muted gains, whereas process-aware criteria such as CoT-pass@k (which require both a correct answer and a syntactically valid chain) often reveal larger deltas in settings explicitly optimized for process rewards (Wen et al., 2025b). We treat these as behavioral signals rather than explanations: chains need not be faithful to internal computation (Chen et al., 2025c). Accordingly, we report both answer- and process-aware metrics alongside calibration (ECE/abstention), and we prefer budget-matched saturation curves over single-point pass@k (section 4).
@@ 3 Perspective I: The RLVR Tax: Benefits, Costs, and Controls
RLVR reliably lifts accuracy on verifiable tasks, but those gains are easy to misread if we ignore systematic costs, the RLVR tax. We elaborate on three recurrent pressures: (i) hallucination and overconfidence, (ii) erosion of instruction following, and (iii) safety and privacy exposure. Each maps directly to our position: without tax-aware controls and sober measurement, reported “reasoning gains” are easy to overstate and hard to transfer.
@@ 3.1 Hallucination and Overconfidence
We make this concrete with a factual-QA control that tracks Not attempted (lower is better), Accuracy (shared) (accuracy on items that both models at a given scale attempted, higher is better), and Expected Calibration Error (ECE; lower is better) computed on shared items from the model’s stated confidence. The dominant effect is attempt inflation: stronger RLVR policies attempt far more questions. For example, moving from DeepSeek-V3 to DeepSeek-R1 (Guo et al., 2025) sharply reduces abstentions and improves shared ECE, but accuracy on the newly attempted tail is modest, so reported gains depend on how one aggregates across attempted vs. unattempted items (Zeng et al., 2025). In other words, scaling RL encourages attempts; whether this yields net quality improvements depends on calibration and gating.
@@ 3.2 Instruction Following
Reasoning-optimized training can erode controllability, especially for long generations. Several studies observe regressions in instruction fidelity when training emphasizes extended chain-of-thought or purely verifiable endpoints (Fu et al., 2025; Li et al., 2025). Moreover, recent evidence shows that models tend to overfit to a small set of verifiable constraints and struggle to generalize; IFBench (58 out-of-domain constraints) documents this gap and finds that RLVR improves precise instruction-following generalization (Pyatkin et al., 2025). This is why our minimum standard in section 6 assumes either an explicit instruction or format component in the reward, or at least a separate instruction-following evaluation pack.
@@ 3.3 Safety and Privacy
Long, explicit reasoning traces widen the attack and leakage surface. Frontier reasoning models are jailbreakable at high rates; automated attacks approach deterministic success in the studied settings, and success rises with attempt and context budgets (Kassianik and Karbasi, 2025; Zhou et al., 2025a; Jiang et al., 2025). Many-shot evaluations show the same scaling. Targeted defenses tuned specifically to that regime can sharply reduce measured attack success in controlled tests, but these gains are brittle across models and budgets and can impose utility costs (Ackerman and Panickssery, 2025). Richer CoT also increases privacy risk: attributes, confidential prompts, and dataset contents can be reconstructed more easily as intermediate reasoning grows (Green et al., 2025). Safety tuning itself can impose a “safety tax,” degrading math and coding unless staged and balanced carefully (Huang et al., 2025; Zhang et al., 2025b).
@@
These taxes distort reported gains predictably: overconfident hallucinations inflate apparent utility, instruction drift undermines deployability, and expanded attack surfaces raise real-world risk. The answer is not to abandon RLVR but to co-optimize for reliability: compose rewards so correctness, grounding, and abstention can all improve; manage variance and difficulty; and make calibration and provenance part of evaluation. We follow this recipe in the next sections to show where conclusions change once the tax is accounted for in section 4 and section 6.
@@ 4.1 Budget Parity and Saturation
A recurring pattern in the literature is to report pass@k for RLVR models while holding base models to much smaller budgets (e.g., pass@1 or pass@5). In such settings, measured improvements often reflect extra search rather than a better policy. Prior work stresses matched budgets and decontaminated baselines for fair comparison (Wen et al., 2025a). Small or underspecified benchmarks inflate estimator variance and reduce statistical power, making single-run point estimates unstable; averaging across multiple randomized trials materially improves estimate reliability (Mu et al., 2025). Under parity controls (e.g., SoberScore, a standardized, matched-budget, multi-seed evaluation that fixes decoding/setup and reports mean $\pm$ std), several prominent gaps shrink or disappear (Hochlehnert et al., 2025). In practice, evaluations should match $k$ across base and RLVR, plot saturation curves (accuracy vs. $k$, optionally summarized by area under the curve), and disclose decoding budgets and parameters.

### 25 May 2026 v3

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) is a practical, scalable way to improve large language models on math, code, and other structured tasks. However, we argue that many headline RLVR gains are not yet well validated because reports often conflate policy improvement with three confounds: (i) budget mismatch between RLVR and baseline evaluations, (ii) attempt inflation and calibration drift that convert abstentions into confident answers, and (iii) benchmark data contamination. Using budget-matched reproductions and partial-prompt contamination probes, we find that several widely cited gaps shrink substantially or disappear once budgets, prompts, and dataset versions are matched and contaminated sets are treated as memorization probes rather than evidence of reasoning. This does not mean that RLVR is ineffective, but it implies that current measurements often overstate capability gains and obscure reliability costs. We therefore propose a compact, tax-aware minimum standard for RLVR training and evaluation: budget-matched saturation curves with variance, calibration, and abstention tracking, a judge-robustness stress test when LLM judges are used, and an explicit contamination screen. With these controls, RLVR remains effective and deployable in verifiable domains, but reasoning gains should be treated as provisional without them.
@@
Fang Wu1*†   Aaron Tu2*   Weihao Xuan3,4*   Heli Qi4,5*   Xu Huang6   Qingcheng Zeng7 Shayan Talaei1   Yijia Xiao8   Peng Xia9   Xiangru Tang10   Yuchen Zhuang6   Yinxi Li11 Bing Hu12   Hanqun Cao13   Wenqi Shi14   Rui Yang15   Nan Liu15   Huaxiu Yao9 Ge Liu16   Li Erran Li17   Amin Saberi1   Naoto Yokoya3,4   Jure Leskovec1   Yejin Choi1† 1Stanford University  2UC Berkeley  3The University of Tokyo  4RIKEN AIP 5Waseda University  6Georgia Tech  7Northwestern University  8UCLA 9UNC Chapel Hill  10Yale University  11University of Waterloo 12Independent Researcher  13CUHK  14UT Southwestern Medical Center 15National University of Singapore  16UIUC  17Amazon AWS AI *Equal Contribution   Corresponding Authors
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a leading post-training route for improving large language models (LLMs) on math and code tasks (Trung et al., 2024; Wen et al., 2025a). By optimizing against automatically computable signals such as unit tests for programs, exact numeric or string matches for math, or retrieval-grounded checks for citations, RLVR promises a path to better performance. Recent work reports large benchmark gains across domains. Appendix Fig. 2 shows a trend plot of RLVR-tagged paper activity and AIME24/25 performance with striking monthly gains. Our position is that RLVR is effective and deployable for verifiable domains, but “reasoning” gains should be treated as provisional without a small core of tax-aware controls.
@@
A central question frames the debate: Does RLVR genuinely impart new reasoning capability, or does it mainly sharpen selection among behaviors the base model already knows how to produce? Budget-parity controlled studies show that base models can narrow or erase RLVR gaps when given matched sampling budgets, consistent with improved search rather than capability expansion (Yue et al., 2025; Wu et al., 2025c). Meanwhile, several settings report gains that are hard to recover by sampling alone: explicitly optimizing the multi-sample objective (pass@k training) (Chen et al., 2025d), curriculum via self-play with variational problem synthesis (Liang et al., 2026), and distribution-aware reward shaping that counters rank bias and diversity collapse, alongside longer-horizon RL schedules and unlikeliness rewards (Liu et al., 2025a; He et al., 2025a). Metric choice can flip conclusions: answer-only scores may diverge from process-aware metrics that require both a correct answer and a valid chain of thought.
@@
A second complication is what we call the RLVR tax: unintended empirical side effects that accompany apparent gains under current reasoning-style post-training (reduced abstention, miscalibration, instruction-fidelity drift, and a larger safety/privacy surface due to longer traces). This tax is neither inevitable nor unique to RLVR; similar patterns arise under reasoning-heavy SFT and RLHF (Wu et al., 2025b). We focus on RLVR because verifiable objectives and open-weight models make these trade-offs measurable at relatively low cost. In practice, RLVR tends to reduce abstention and increase stated confidence, sometimes even when answers are wrong, thereby shifting risk from “I don’t know” to assertive errors (Song et al., 2025; Mei et al., 2025; Yao et al., 2025). It can also erode instruction fidelity on longer generations, where adhering to formats or constraints becomes harder as chains grow (Fu et al., 2025; Li et al., 2025). Finally, longer, more explicit reasoning traces expand the attack and leakage surface, raising jailbreak success and privacy exposure if left unchecked (Zhou et al., 2025a; Jiang et al., 2025; Ackerman and Panickssery, 2025; Green et al., 2025; Huang et al., 2025; Zhang et al., 2025).
@@
A third complication is measurement. Reported advances are sensitive to sampling budgets and decoding settings (Hochlehnert et al., 2025; Chandak et al., 2025; Brown et al., 2024; Muennighoff et al., 2025), to the stability of LLM-as-a-judge evaluators (Zhao et al., 2025b; Tan et al., 2024; Gu et al., 2024), to calibration drift (Leng et al., 2024; Xiao et al., 2025), and to data provenance (He et al., 2025b; Mirzadeh et al., 2024; Jain et al., 2024; Xu et al., 2025; Wu and Li, 2024; Wu et al., 2025a, 2026a). When budgets are matched, judges are stress-tested, and datasets are versioned and decontaminated, reported improvement gaps can diminish, suggesting that part of the apparent progress reflects evaluation design rather than durable capability (Hochlehnert et al., 2025; Chandak et al., 2025). We organize the paper around three threads: the RLVR tax, evaluation pitfalls, and data contamination, culminating in a compact tax-aware minimum standard. A visual roadmap is in Fig. 1.
@@
We use RLVR to mean post-training that optimizes LLMs against automatically checkable signals rather than human preference models (Ouyang et al., 2022). Typical implementations adopt PPO/GRPO/DAPO-style updates with a KL penalty to the base policy, often mixing offline rollouts with online sampling, adding entropy regularization, and using selective filtering or gating. Rewards are frequently componentized (for correctness, grounding/citation sufficiency, and calibrated refusal) and introduced in stages. Our analysis centers on open-weight models fine-tuned for math, code, and question answering (QA) with verifiable objectives, in both single- and multi-domain regimes. We deliberately avoid a catalogue of low-level algorithmic variants except where they bear on reliability (e.g., unlikeliness rewards that counter rank bias, or curricula that change variance). The primary threats considered are privacy leakage from long chains of thought (CoT) and increased jailbreak susceptibility during or after RLVR.
@@
Our contributions are threefold: (1) Measurement: we provide budget-matched reproductions and a unified gap table that isolates the effects of budget mismatch, judge drift, template drift, and dataset versioning (Table 2). (2) Tax: we quantify attempt inflation and the resulting calibration and refusal costs under RLVR and matched SFT controls (Table 1). (3) Protocol: based on these findings, we distill a compact tax-aware minimum standard for RLVR training and evaluation, intended as a minimum credible reporting bar, which we instantiate in section 6. The paper is organized accordingly: we (i) synthesize evidence on sharpening vs. expansion (section 2); (ii) analyze the RLVR tax and how it distorts reported improvements (section 3); (iii) reproduce representative gaps under parity controls (section 4); (iv) audit contamination with partial-prompt probes (section 5); and (v) present the tax-aware minimum standard (section 6).
@@ 2.1 Skeptical Lens: Sharpening and Sampling Effects
RLVR appears to improve sample efficiency and steer the model toward high-reward regions already present in the base distribution, rather than broaden fundamental capability (Gandhi et al., 2025; Shah et al., 2025). When base models are evaluated under matched sampling budgets (i.e., pass@k), gaps between base and RLVR-trained models often shrink (Yue et al., 2025), consistent with sharpening rather than expansion. Empirically, outputs can collapse toward a dominant pretraining mode with reduced diversity (Zhao et al., 2025a), and theory suggests that after RLVR, support shrinkage tends to outweigh support expansion (Wu et al., 2025c). Stress tests add caution: performance collapses at higher complexities have been reported even with ample compute (Shojaee et al., 2025), although some failures trace to evaluation artifacts such as token limits and unsolvable items, underscoring the need for budget parity, seed control, and dataset hygiene (see section 4). Finally, several works separate knowledge injection from policy optimization: pipelines that use SFT or distillation to inject knowledge and RLVR to select among candidates typically beat pure GRPO, suggesting RLVR optimizes within learned support (Ma et al., 2026; Wan et al., 2025; Chen et al., 2026); guidance and self-distillation further help over vanilla GRPO (Nath et al., 2025; Liu et al., 2025b).
@@ 2.2 Optimistic Lens: Signals of Capability Broadening
Answer-only metrics (e.g., pass@k) can show muted gains, whereas process-aware criteria such as CoT-pass@k (which require both a correct answer and a syntactically valid chain) often reveal larger deltas in settings explicitly optimized for process rewards (Wen et al., 2025b). We treat these as behavioral signals rather than explanations: chains need not be faithful to internal computation (Chen et al., 2025b). Accordingly, we report both answer- and process-aware metrics alongside calibration (ECE/abstention), and we prefer budget-matched saturation curves over single-point pass@k (section 4).
@@ 3 Perspective I: The RLVR Tax: Benefits, Costs, and Controls
RLVR reliably lifts accuracy on verifiable tasks, but those gains are easy to misread if we ignore systematic costs—the RLVR tax. We elaborate on three recurrent pressures: (i) hallucination and overconfidence, (ii) erosion of instruction following, and (iii) safety and privacy exposure. Each maps directly to our position: without tax-aware controls and sober measurement, reported “reasoning gains” are easy to overstate and hard to transfer.
@@ 3.1 Hallucination and Overconfidence
We make this concrete with a factual-QA control that tracks Not attempted (lower is better), Accuracy (shared) (accuracy on items that both models at a given scale attempted; higher is better), and Expected Calibration Error (ECE; lower is better) computed on shared items from the model’s stated confidence. The dominant effect is attempt inflation: stronger RLVR policies attempt far more questions. For example, moving from DeepSeek-V3 to DeepSeek-R1 (DeepSeek-AI, 2025) sharply reduces abstentions and improves shared ECE, but accuracy on the newly attempted tail is modest, so reported gains depend on how one aggregates across attempted and unattempted items (Zeng et al., 2025). In other words, scaling RL encourages attempts; whether this yields net quality improvements depends on calibration and gating.
@@ 3.2 Instruction Following
Reasoning-optimized training can erode controllability, especially for long generations. Several studies observe regressions in instruction fidelity when training emphasizes extended chain-of-thought or purely verifiable endpoints (Fu et al., 2025; Li et al., 2025). At the same time, recent evidence shows that models tend to overfit to a small set of verifiable constraints and struggle to generalize; IFBench (58 out-of-domain constraints) documents this gap and also finds gains under RLVR in precise instruction-following generalization (Pyatkin et al., 2025). This is why our minimum standard in section 6 requires either an explicit instruction or format component in the reward, or at least a separate instruction-following evaluation pack.
@@ 3.3 Safety and Privacy
Long, explicit reasoning traces widen the attack and leakage surface. Frontier reasoning models are jailbreakable at high rates; automated attacks approach deterministic success in the studied settings, and success rises with attempt and context budgets (Kassianik and Karbasi, 2025; Zhou et al., 2025a; Jiang et al., 2025). Many-shot evaluations show the same scaling. Targeted defenses tuned specifically to that regime can sharply reduce measured attack success in controlled tests, but these gains are brittle across models and budgets and can impose utility costs (Ackerman and Panickssery, 2025). Richer CoT also increases privacy risk: attributes, confidential prompts, and dataset contents can be reconstructed more easily as reasoning traces grow longer (Green et al., 2025). Safety tuning itself can impose a “safety tax,” degrading math and coding unless staged and balanced carefully (Huang et al., 2025; Zhang et al., 2025).
@@
These taxes distort reported gains predictably: overconfident hallucinations inflate apparent utility, instruction drift undermines deployability, and expanded attack surfaces raise real-world risk. The answer is not to abandon RLVR but to co-optimize for reliability: compose rewards so correctness, grounding, and calibrated abstention can all improve; manage variance and difficulty; and make calibration and provenance part of evaluation. We follow this recipe in the next sections to show where conclusions change once the tax is accounted for in section 4 and section 6.
@@ 4.1 Budget Parity and Saturation
A recurring pattern in the literature is to report pass@k for RLVR models while holding base models to much smaller budgets (e.g., pass@1 or pass@5). In such settings, measured improvements often reflect extra search rather than a better policy. Prior work stresses matched budgets and decontaminated baselines for fair comparison (Hochlehnert et al., 2025; Chandak et al., 2025; Yue et al., 2025). Small or underspecified benchmarks inflate estimator variance and reduce statistical power, making single-run point estimates unstable; averaging across multiple randomized trials materially improves estimate reliability (Mu et al., 2025). Under parity controls (e.g., SoberScore, a standardized, matched-budget, multi-seed evaluation that fixes decoding/setup and reports mean $\pm$ std), several prominent gaps shrink or disappear (Hochlehnert et al., 2025). In practice, evaluations should match $k$ across base and RLVR, plot saturation curves (accuracy vs. $k$, optionally summarized by area under the curve), and disclose decoding budgets and parameters.